OpenStudy (anonymous):

Which formula would you use when you are looking for the slope of a line and given two points on the line?

OpenStudy (anonymous):

line? what line

OpenStudy (anonymous):

??

OpenStudy (anonymous):

A. point-slope formula B. slope formula C. slope formula and point-slope formula D. slope-intercept formula

OpenStudy (anonymous):

c

Nnesha (nnesha):

slope intercept form\[y = mx +b \] point slope form \[y -y _{1} =m (x -x _{1})\] slope formula \[m = \frac{ y _{2}-y _{1} }{ x _{2} - x _{1}}\]